• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 00:03:25 on February21,2002

    Entry number 83307

    keyword=Swing shift summary

    16:00 Begin shift in maintenance. Left spectrometer momentum set at 3.1760 GeV/c and angle at 18.43 degrees.
    16:20 Trigger downloaded for Kinematics 3E, and calorimeter threshold set at 370.

    16:23-19:31 Run # 2717 cosmics, Carbon target
    17:10 Request MCC to put Hall A in Power Permit
    18:15 Called MCC, and they said that they may deliver beam in less than an hour. Also requested they put us in Beam Permit
    19:00 Called MCC, and we may have beam in about 30 minutes.
    19:00 Recycled Q2 and Q3 for Left Spectrometer.
    19:36-20:00 Run # 2718 cosmics, Carbon target
    19:37 Hall A in Beam Permit.
    19:40 MCC called. They are cycling the magnets and will send tuned beam soon.
    19:52 MCC is preparing to send use CW beam. Requested 20 uA, unrasted.
    20:05 Beam position off. Asked for to correct for x = 0.05 and y = 1.70

    20:09 Asked controlled Access to check why T5 is zero.
    20:34 Trigger fixed for T5. Hall A going to Power Permit. Cable wasn't connected.
    20:37 Asked for Beam Permit and 20 uA unrastered beam.
    20:40 Hall A in Beam Permit
    20:46 Sending beam to Hall A
    20:47-21:08 Run # 2719, Carbon, moff, Veto off
    21:08 Mask Target to move to LH2 non-radiator
    21:12 Target in place. Asked for 40 uA, rastered beam.
    21:13-21:25 Run # 2720, LH2 non-rad, moff, Veto off
    21:25 Turn calorimeter magnet on.
    21:27-21:38 Run # 2721, LH2 non-rad, mon, Veto off
    21:42 Magnet off
    21:42-21:53 Run # 2722, LH2 non-rad, moff, Veto off
    21:53-22:05 Run # 2723, LH2 non-rad, moff, Veto off
    22:05-22:18 Run # 2724, LH2 non-rad, moff, Veto off

    22:19 Ask MCC to mask target to move to LH2 + rad and to change spectrometer.
    22:28 Target in LH2 + rad position. Left spectrometer anlge at 19.47 degrees. Ask MCC for 15 uA of beam with raster on.
    22:33 Left spectrometer momentum set at 2.939 GeV/c and calorimeter magnet on.
    22:33-23:08 Run #'s 2725 - 2729, LH2 + rad, mon, Veto off
    23:10 Change calorimeter threshold to 420
    23:18 Run # 2730 Pedestal Run, same conditions as before
    23:20-23:47 Run # 2731, LH2 + rad, mon, Veto off
    23:51-? Run # 2732, LH2 + rad, mon, Veto off

    00:00 Owl shift takes over with 2732 still running.